OpenStudy (anonymous):

Solve the equation 1296 x = 6

OpenStudy (anonymous):

\[1296x=6\]\[x=\frac{6}{1296}\]\[x=\frac{1}{216}\]You should try putting more effort into this.

OpenStudy (anonymous):

Divide both sides of the equation by 1296 to isolate x. 1296x/1296 = 6/1296 x = 6/1296 Which simplifies to x = 1/216

OpenStudy (saifoo.khan):

Right^

OpenStudy (anonymous):

I'm not good at math

OpenStudy (anonymous):

 LIKE FOR GOOD ANSWER